# Routine haematological markers can predict and discriminate health status and biological age even from noisy sources

### v10

@@ Abstract
For more than two decades, advances in personalised medicine and precision healthcare have largely been based on genomics and other omics data. These strategies aim to tailor interventions to individual patient profiles, promising greater treatment efficacy and more efficient allocation of healthcare resources. Here, we show that widely collected common haematological markers can reliably predict and discriminate individual chronological age and health status from even noisy sources. Our analysis includes synthetic and real retrospective patient data, including medically relevant and extreme cases. We combine fully explainable risk assessment scores with machine and deep learning techniques to focus on clinically significant patterns and characteristics without functioning purely as a “black box” model allowing interpretation and control. We validated the results with the CDC NHANES database (100 000 participants) and the UK Biobank (500 000 participants), the former a noisy survey and the second a noisy longitudinal study. Unlike current biological ageing indicators, this approach may offer rapid, and scalable implementations of personalised, precision and predictive approaches to healthcare and medicine without or before requiring other specialised, uncommon or costly tests.
@@ 1 Introduction
In a recent retrospective study, Complete Blood Count tests of 12 000 adults over 20 years showed that individuals are stable over time while being healthy and have unique characteristics that effectively act like health fingerprints, specific patterns or personalised signatures that when deviate from their personal baselines are indicative of health and even mortality [33]. These personalised benchmarks can significantly improve early disease detection by providing a better and more precise picture of individual health. This had already been previously suggested from more specialised immunological tests [27, 28]. Unlike the study [33], the present work focusses on defining a minimal, interpretable index or risk assessment score grounded in routine haematological and circulatory biomarkers. Taking advantage of population-normalised Complete Blood Count (CBC) parameters, our approach achieves discriminatory performance not explored in [33] while remaining computationally efficient, reproducible and transparent for clinical purposes. The index therefore is the proactive natural consequence of what was reported in  [33] regarding personal setpoints and clinically relevant deviations even within normal population reference values.
Here, we investigate whether these findings can be leveraged for risk assessment by identifying clinically relevant immune patterns. We further demonstrate that the same test allows robust detection of patterns predictive of age and health status, even in large, retrospective, and noisy datasets. In this paper [1], it was shown that 60 blood markers were predictive of biological age. While our paper precedes theirs as posted online, in this one we show that only a small subset of those markers are actually necessary for predictive purposes and that ad hoc artificial weighting of markers is not necessary.
@@
Building on this need, we propose immune-based indexes derived from standard blood tests as quantitative predictors of health risk and immune ageing, offering valuable tools for preventive and precision medicine. Risk assessment tools are often needed to improve patient safety and patient outcomes. Risk assessment scores used for initial patient classification or diagnosis and triaging are often qualitative and inter-subjective. Robust quantitative tools with clinical relevance are required to optimise health monitoring choices across all levels of healthcare systems. Digital medicine aims to improve the quality of patient care and outcome trajectories through clinical informatics, such as risk assessment metrics grounded in clinical data.
Risk scores are valuable predictive tools that, when implemented, can be used by clinicians with positive results, leading to better patient outcomes in precision medicine. Risk assessment measures for patient triage are often needed to optimise patient pathways and health monitoring across all levels of healthcare. However, most scores are defined for very specific conditions, such as heart disease, cardiovascular disease (CVD) or venous thromboembolism (VTE), to name just a few. Others reflect factors such as risk of severe disease [5, 6], clinical acuity [7, 8], and long-term outcomes [9, 10]. For example, the Intermountain risk index [12] provides a Complete Blood Count (CBC/FBC) risk index based on the size of red blood cells, shown to be associated with bleeding and in-hospital mortality. Originally used as an anaemia predictor [11], it is often reported in FBC/CBC panels. As precision healthcare and digital medicine continue to advance, these risk scores will likely integrate more advanced data-driven approaches to improve prediction accuracy and ensure relevance across a broad spectrum of clinical conditions.
@@ 1.1 Biological Ageing
Biological ageing clocks have been explored from various perspectives and approaches from skin to epigenetic [2] using blood to microbime [3] (skin and gut) to proteomic RNA [4]. Epigenetic measurements are limited in scope due to limited range of molecules measures while proteomics offers a comprehensive means for measuring ageing but whole-body proteomic signals are noisy due to different organ contribution to the full signal and skin ageing for proteomics leads to tissue damage while gut microbiome and all others require specialised not routine tests.
Immune age-related scores based on inflammation markers have also traditionally required highly specialised and therefore uncommon and expensive tests that are not suitable for mass adoption or mass deployment in current clinical pathways, as they are difficult to scale to cover large populations [14, 32, 15, 18, 16].
@@
Routine blood tests have been associated with the prediction of short-term mortality and with the ability to significantly improve emergency department triage [46]. Most risk scores depend on multiple external factors that are unique to a single test result in and of itself, and have proven useful when available, even in the context of the immune system [31]. Several of these scores are based on factors such as physiological data, medical and family history, and lifestyle choices, and reference a range of things, from age and sex to drinking and smoking habits. Although numerical scores and colour-based scores are relatively common, no single general numerical and colour-based blood test risk assessment index exists that is based on a combination of distances from the population and personalised average values, in particular for CBC/FBC and immune-related data, used to guide Machine Learning approaches for discrimination of individual health predictors.
@@ 2 Methods
A Complete Blood Count (CBC) or Full Blood Count (FBC) test is the most popular blood test with over 4 Billion of them performed every year world wide, 500 million only in the U.S. It traditionally includes the total number of red blood cells (RBCs); the total number of white blood cells (WBCs) and platelets (PLT); haemoglobin (HGB); haematocrit (HCT); estimations related to concentration, shape and volume of red cells (MCV, RDW, MCH and MCHC) and counts of different WBC subtypes, including Lymphocytes (LYMP), Monocytes (MONO), Neutrophils (NEUT), Eosinophils (EOS) and Basophils (BAS). A CBC is a routine blood test used to assess overall health and detect a wide range of disorders, including anaemia, infection and leukaemia.
The CBC test is deeply related to the immune system and is a key diagnostic tool that can provide information on a variety of diseases and conditions, including infection, inflammation, cancer, anaemia, heart disease, and more. CDC NHANES data were used to estimate deficiencies and toxicities of specific nutrients in the population and subgroups in the US, to extract population reference data, and to estimate the contribution of diet, supplements, and other factors to whole blood levels of nutrients. This repository includes CBC data can be used for research purposes and is publicly available [17].
@@ 2.1 Numerical index and Age Prediction
NIS (Numerical Immune Score) Quantitative index derived from haematological analyte deviations; represents raw numerical immune deviation without categorical scaling.
CCIS (Colour-Coded Immune Score) Categorical extension of NIS using colour-coded bins to improve interpretability and visual comparison between individuals.
NCCIS (Normalised Colour-Coded Immune Score) Normalised version of CCIS adjusted to population reference intervals (NHANES, NHS-UK), allowing generalisation and comparison across diverse datasets and laboratories.
@@
The sex groups were separated by health status determined by dividing immune scores and ages by the healthy range thresholds indicated in Table  S1 (see Sup. Inf.), which are as follows: Red Blood Cell (RBC) count (between 3.8 to 5.8), and the following healthy ranges for the multivariate measure: normal Mean Corpuscular Haemoglobin (MCH) (27-33), Haematocrit (HCT) (37-47), lymphocytes (LYMP) (1.0-4.0), and neutrophils (NEUT) (1.8-7.5). Above or below these healthy count ranges, the individuals were categorised into unhealthy groups. In addition, error rates analysis for age prediction between actual ages and immune index computed ages was performed using these binned subsets.
The zero counts were removed from the NHANES data prior to all analysis. Zero values were excluded only when they represented non-physiological or missing data artifacts (e.g., unrecorded analytes). True physiological zeros, where relevant, were retained. However, for the quantitative analysis, sex was replaced by 1 for males and 2 to denote females. Table S2 (see Sup. Inf.) highlights sex-specific differences in immune analyte values, for reference. Because the NHANES and UK Biobank datasets compile data from multiple laboratories, we standardized all analyte values using the corresponding national population reference intervals (NHANES and NHS-UK) prior to computing the indices. This population-level normalization mitigates inter-laboratory cut-off variability and supports cross-site comparability.
@@
Ageing is associated with complex changes and dysregulation of cellular processes. Nine tentative hallmarks that represent common denominators of ageing in different organisms have been described [42].
As lifespan increases in many countries, there is a concomitant increase in age-related morbidities, particularly cardiovascular disease, and increased susceptibility to seasonal infections, cancers, and neurodegenerative disorders. A unifying link between the higher rates of these disparate diseases observed in ageing populations is the progressive decline in immune function.
It has been postulated that inflammation plays a critical role in the regulation of physiological ageing. Inflammatory components of the immune system are often chronically elevated in aged individuals, a phenomenon that has been termed “inflammageing” [34, 35, 36, 37, 38, 18].
However, the dynamics of this process at the individual level have not been characterised, hindering quantification of an individual’s ‘immune age’. In Ref. [18], various ‘omics’ technologies are used to capture population- and individual-level changes in the human immune system in individuals of different ages sampled longitudinally over a nine-year period. They observed high inter-individual variability in the rates of change of cellular frequencies that was dictated by the individual’s baseline values, allowing identification of steady-state levels toward which a cell subset converged and the ordered convergence of multiple cell subsets toward older adult homeostasis. In [39], an IMM-AGE index was described that captured an individual’s immune-ageing process. The IMM-AGE index was correlated with age, yet it captured additional metrics such as cell-cytokine response better than chronological age.
@@
The UK Biobank is a long-term prospective nationwide study that houses the deidentified biological samples and health-related data of half a million people. Between 2006 and 2010 volunteers aged 40 to 69 years were recruited from all over Great Britain and consented to share their health data and to be followed for at least 30 years thereafter; with the objective of allowing scientific discoveries to be made in the prevention, diagnosis, and treatment of disease [21, 23]. We tested the NIS index (which is the NCCIS index without the warning colouring component, see Supp. Inf.), to separate large disease categories in the UK Biobank. The database does not contain a clean healthy control group, meaning that all participants had potentially some reported health problem, making our separation results more relevant given that we had to create a Rest group different from the large disease categories identified as a noisy control group.
@@ 3 Results
Figure 1: Distribution of the Normalised Colour-Coded Immune index (NCCIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000-2018 database. Individuals are defined as Healthy if all their analytes are within the normal ranges as defined by the NHS (S1 (see Sup. Inf.) and Unhealthy otherwise. Self rep. healthy comprises individuals who self-reported health of 3 (“Good”) or less. HIV is an example of chronic infection while all the diseases can fall in the category of immune and blood-related conditions.While separation between overtly healthy and disease groups may be expected, the novelty of our proposed metric lies in its ability to provide a generalizable, quantitative, and interpretable single numerical score that scales across heterogeneous conditions and noisy datasets, offering a standardized baseline for predictive stratification.
Table 5 shows examples of CCIS and its corresponding NCCIS values for a sample of real-life and artificial cases. The second column of the table shows the NIS, for purposes of comparison. It is worth noting that these examples were generated using the NHS values for normal healthy subjects and based on the provisional assumption that the distance between the upper and normal ranges is twice the standard deviation. The examples were taken from [22] as they were intended as very preliminary tests. We are aware that their source clearly states that they are meant only for teaching purposes. Here, they are used for purposes of illustration only, and the development of the NIS and NCCIS is not dependent on them. In Table 4, we provide Cohen’s d values that suggest practical applicability of the index, making it clinically relevant for triaging purposes even at the level of individual classification and prediction.
@@
The method is shown to be valid for predicting immune profiles/trends (variation over time) observed in diverse patient groups when considering socio-clinical metadata parameters such as age and sex, thereby suggesting their clinical applicability for purposes such as triaging and fast screening and extensions to deviant clinical cases in prospective studies. In Fig. 1, it is shown that reported medical conditions produced a higher average NCCIS value, but also induced some anomalous clustering. Furthermore, the distribution of scores was even wider than that of the NIS. We therefore conclude that the NCCIS did enhance the insights already gleaned from the NIS and the CCIS.Separation of groups of self-reported data suggests that they could be separated from those that were not actually healthy at the time of testing. The self-reported health information may not always correspond to objectively verified clinical status. Thus, the observation possibly reflects statistical discrepancies between self-reported and haematological patterns rather than confirmed diagnostic misclassification. Accordingly, while individuals who self-reported not being healthy generally exhibited blood profiles consistent with non-healthy patterns, this is suggested as a potential interpretation rather than clinical distinction. The number of individuals in each cohort is shown in Table 3. One possible explanation for the bimodality of some of these plots is the nature of the source data as self-reported, where respondents may or may not have been ill at the time of giving their answers, or from the application of the separating cutoff value from NIS.
@@
Statistical tests assessed the relationship among the three individualised scores, NIS, NCCIS, and CCIS, for various health conditions. Table 6 presents significant statistical associations between these scores. The Kruskal-Wallis test, a non-parametric alternative to ANOVA was performed due to the ordinal nature of the CCIS index. The significant p-value (<<0.001) indicates that there are statistically significant differences in the distribution of the NIS, NCCIS, and CCIS scores in the conditions. A Pearson correlation assessed the linear relationship between NIS and NCCIS. The strong positive correlation coefficient (0.89) with a highly significant p-value (<< 0.0001) determined from a two-tailed independent samples t-test suggests a robust linear association between these two scores. Lastly, a Wilcoxon Signed-Rank test compared paired differences between NIS and NCCIS. The significant result (p = 0.018) indicates that there are significant differences between these paired scores and therefore each can capture different aspects of blood data.

### v12

# Integrative adaptive indexes from noisy routine haematological markers can predict and discriminate health status and biological age

@@ Abstract
For more than two decades, advances in personalised medicine and precision healthcare have largely been based on genomics and other omics data. These strategies aim to tailor interventions to individual patient profiles, promising greater treatment efficacy and more efficient allocation of healthcare resources. Here, we show that widely collected common haematological markers can reliably predict and discriminate individual chronological age and health status from even noisy sources. Our analysis includes synthetic and real retrospective patient data, including medically relevant and extreme cases. We combine fully explainable risk assessment scores with machine intelligence to focus on clinically significant patterns and characteristics without functioning as a “black box” model and allowing interpretation and control tested on the US CDC NHANES database (100 000 participants) and validated with the UK Biobank (500 000 participants). Despite the noisy nature of these databases due to self-reporting and sparse data, the results remained relevant and statistically significant. Unlike current biological ageing indicators, this approach may offer rapid, and scalable implementations for personalised, precision and predictive approaches to healthcare and medicine without or before requiring other specialised, uncommon, or costly tests.
@@ 1 Introduction
In a recent retrospective study, Complete Blood Count tests of 12,000 adults over 20 years showed that individuals are stable over time while being healthy and have unique characteristics that effectively act like health fingerprints, specific patterns or personalised signatures that when deviate from their personal baselines are indicative of health and even mortality [40].
These personalised benchmarks can significantly improve early disease detection by providing a better and more precise picture of individual health as suggested in various immunological studies [34, 35]. In contrast to previous studies [40], the present work focusses on defining a minimal, interpretable index or risk assessment score grounded in routine haematological and circulatory biomarkers. Taking advantage of population-normalised Complete Blood Count (CBC) parameters, our approach achieves discriminatory performance not explored in [40] while remaining computationally efficient, reproducible and transparent for clinical purposes. The index therefore is the proactive natural consequence of what was reported in  [40] regarding personal setpoints and clinically relevant deviations even within normal population reference values.
We investigate whether these findings can be leveraged for risk assessment by identifying clinically relevant immune patterns. We further demonstrate that the same test allows robust detection of patterns predictive of age and health status, even in large, retrospective, and noisy datasets. In this paper [1], it was shown that 60 blood markers were predictive of biological age. While our paper precedes theirs as posted online, in this one we show that only a small subset of those markers are actually necessary for predictive purposes and that ad hoc artificial weighting of markers is not necessary.
@@
Building on this need, we propose immune-based indexes derived from standard blood tests as quantitative predictors of health risk and immune ageing, offering valuable tools for preventive and precision medicine. Risk assessment tools are often needed to improve patient safety and patient outcomes.
Risk assessment scores used for initial patient classification or diagnosis and triaging are often qualitative and inter-subjective. Robust quantitative tools with clinical relevance are required to optimise health monitoring choices across all levels of healthcare systems. Digital medicine aims to improve the quality of patient care and outcome trajectories through clinical informatics, such as risk assessment metrics based on clinical data.
Risk scores are valuable predictive tools that, when implemented, can be used by clinicians with positive results, leading to better patient outcomes in precision medicine. Risk assessment measures for patient triage are often needed to optimise patient pathways and health monitoring across all levels of healthcare.
However, most scores are defined for very specific conditions, such as heart disease, cardiovascular disease (CVD) or venous thromboembolism (VTE), to name just a few. Others reflect factors such as risk of severe disease [12, 13], clinical acuity [14, 15], and long-term outcomes [16, 17].
For example, the Intermountain risk index [19] provides a Complete Blood Count (CBC/FBC) risk index based on the size of red blood cells, shown to be associated with bleeding and in-hospital mortality. Originally used as an anaemia predictor [18], it is often reported in FBC/CBC panels.
As precision healthcare and digital medicine continue to advance, these risk scores will likely integrate more advanced data-driven approaches to improve prediction accuracy and ensure relevance across a broad spectrum of clinical conditions.
@@ 1.1 Biological Ageing
Biological ageing clocks have been explored from various perspectives and approaches from skin to epigenetic [2] using blood to microbime [3] (skin and gut) to proteomic RNA [4].
Epigenetic measurements are limited in scope due to limited range of molecules measures, while proteomics offers a comprehensive means for measuring ageing. However, whole-body proteomic signals are noisy due to different organ contribution to the full signal, and skin ageing for proteomics leads to tissue damage.
Gut microbiome and other age-related scores, require specialised tests and not routine tests. For example, immune age-related scores based on inflammation markers have traditionally required highly specialised and therefore uncommon and expensive tests that are not suitable for mass adoption or mass deployment in current clinical pathways, as they are difficult to scale to cover large populations [21, 39, 22, 25, 23].
@@
Routine blood tests have been associated with the prediction of short-term mortality and with the ability to significantly improve emergency department triage [53].
Most risk scores depend on multiple external factors that are unique to a single test result in and of itself, and have proven useful when available, even in the context of the immune system [38]. Several of these scores are based on factors such as physiological data, medical and family history, and lifestyle choices, and reference a range of things, from age and sex to drinking and smoking habits.
Although numerical scores and colour-based scores are relatively common, no single general numerical and colour-based blood test risk assessment index exists that is based on a combination of distances from the population and personalised average values, in particular for CBC/FBC and immune-related data, used to guide Machine Learning approaches for discrimination of individual health predictors.
@@
Further, from a mechanistic viewpoint, age-associated perturbations or changes in CBC analytes may reflect early detection markers of ageing-related disease progression and immunosenescence processes, including myeloid skewing, reduced lymphopoiesis, increased inflammatory tone (“inflammaging”), clonal hematopoiesis (or of intermediate potential) and epigenetic or genetic alterations in hematopoietic stem cell population dynamics, erythropoiesis, and platelet activation  [21, 39, 5, 6, 7]. Chronic psychosocial and physiological stressors have also been linked to sustained immune activation and epigenetic remodeling of hematopoietic lineages, potentially accelerating immune-aging phenotypes detectable in peripheral blood indices [8, 9]. Emerging discussions in the brain-gut-immune axes are key converging points. CBC indices have been independently associated with systemic inflammatory burden, immune-microenvironmental remodelling, epigenetic alterations, and vascular ageing in prior literature. [10, 11].
Thus, the composite index integrates distributed but biologically grounded ageing-related predictive signatures.
@@ 2 Methods
A Complete Blood Count (CBC) or Full Blood Count (FBC) test is the most popular blood test with over 4 Billion of them performed every year world wide, 500 million only in the U.S.
It traditionally includes the total number of red blood cells (RBCs); the total number of white blood cells (WBCs) and platelets (PLT); haemoglobin (HGB); haematocrit (HCT); estimations related to concentration, shape and volume of red cells (MCV, RDW, MCH and MCHC) and counts of different WBC subtypes, including Lymphocytes (LYMP), Monocytes (MONO), Neutrophils (NEUT), Eosinophils (EOS) and Basophils (BAS).
A CBC is a routine blood test used to assess overall health and detect a wide range of disorders, including anaemia, infection and leukaemia. A CBC test is deeply related to the immune system and is a key diagnostic tool that can provide information on a variety of diseases and conditions, including infection, inflammation, cancer, anaemia, heart disease, and more. CDC NHANES data were used to estimate deficiencies and toxicities of specific nutrients in the population and subgroups in the US, to extract population reference data, and to estimate the contribution of diet, supplements, and other factors to whole blood levels of nutrients. This repository includes CBC data can be used for research purposes and is publicly available [24].
@@ 2.1 Numerical index and Age Prediction
While CBC measurements may exhibit slight calibration differences across institutions, both NHANES and the UK Biobank inherently incorporate multi-site variability. Thereby, the data embeds real-world technical heterogeneity into the training distributions. Population-level normalization lowers systematic variations, and the adaptive personalised baseline further reduces sensitivity to fixed calibration bias. In prospects, if laboratory-specific calibration coefficients are available, the framework can incorporate them as buffers or corrections prior to index computation, preserving generalisability while maintaining clinical interpretability.
@@
NIS
(Numerical Immune Score) Quantitative index derived from haematological analyte deviations; represents raw numerical immune deviation without categorical scaling.
CCIS
(Colour-Coded Immune Score) Categorical extension of NIS using colour-coded bins to improve interpretability and visual comparison between individuals.
NCCIS
(Normalised Colour-Coded Immune Score) Normalised version of CCIS adjusted to population reference intervals (NHANES, NHS-UK), allowing generalisation and comparison across diverse datasets and laboratories.
@@
The sex groups were separated by health status determined by dividing immune scores and ages by the healthy range thresholds indicated in Table  S1 (see Sup. Inf.).
These thresholds are as follows: Red Blood Cell (RBC) count (between 3.8 to 5.8), and the following healthy ranges for the multivariate measure: normal Mean Corpuscular Haemoglobin (MCH) (27-33), Haematocrit (HCT) (37-47), lymphocytes (LYMP) (1.0-4.0), and neutrophils (NEUT) (1.8-7.5). Above or below these healthy count ranges, the individuals were categorised into unhealthy groups.
In addition, error rates analysis for age prediction between actual ages and immune index computed ages was performed using these binned subsets. Zero counts were removed from the NHANES data prior to all analysis. Zero values were excluded only when they represented non-physiological or missing data artifacts (e.g., unrecorded analytes). True physiological zeros, where relevant, were retained. However, for the quantitative analysis, sex was replaced by 1 for males and 2 to denote females. Table S2 (see Sup. Inf.) highlights sex-specific differences in immune analyte values, for reference. Because the NHANES and UK Biobank datasets compile data from multiple laboratories, we standardized all analyte values using the corresponding national population reference intervals (NHANES and NHS-UK) prior to computing the indices. This population-level normalization mitigates inter-laboratory cut-off variability and supports cross-site comparability.
@@
Ageing is associated with complex changes and dysregulation of cellular processes. Nine tentative hallmarks that represent common denominators of ageing in different organisms have been described [49]. As lifespan increases in many countries, there is a concomitant increase in age-related morbidities, particularly cardiovascular disease, and increased susceptibility to seasonal infections, cancers, and neurodegenerative disorders. A unifying link between the higher rates of these disparate diseases observed in ageing populations is the progressive decline in immune function.
It has been postulated that inflammation plays a critical role in the regulation of physiological ageing. Inflammatory components of the immune system are often chronically elevated in aged individuals, a phenomenon that has been termed “inflammageing” [41, 42, 43, 44, 45, 25]. However, the dynamics of this process at the individual level have not been characterised, hindering quantification of an individual’s ‘immune age’. In previous work [25], various ‘omics’ technologies are used to capture population- and individual-level changes in the human immune system in individuals of different ages sampled longitudinally over a nine-year period. They observed high inter-individual variability in the rates of change of cellular frequencies that was dictated by the individual’s baseline values. This allowed identification of steady-state levels toward which a cell subset converged and the ordered convergence of multiple cell subsets toward older adult homeostasis.
In [46], an IMM-AGE index was described that captured an individual’s immune-ageing process. The IMM-AGE index was correlated with age, yet it captured additional metrics such as cell-cytokine response better than chronological age.
@@
The UK Biobank is a long-term prospective nationwide study that houses the deidentified biological samples and health-related data of half a million people. Between 2006 and 2010 volunteers aged 40 to 69 years were recruited from all over Great Britain and consented to share their health data and to be followed for at least 30 years thereafter; with the objective of allowing scientific discoveries to be made in the prevention, diagnosis, and treatment of disease [28, 30].
We tested the NIS index (which is the NCCIS index without the warning colouring component, see Supp. Inf.), to separate large disease categories in the UK Biobank.
The database does not contain a clean healthy control group, meaning that all participants had potentially some reported health problem, making our separation results more relevant given that we had to create a Rest group different from the large disease categories identified as a noisy control group.
@@ 3 Results
Figure 1: Distribution of the Normalised Colour-Coded Immune index (NCCIS) and Numerical Immune Score (NIS).
Figure 2: Distribution of the Normalised Colour-Coded Immune index (NCCIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000-2018 database. Individuals are defined as Healthy if all their analytes are within the normal ranges as defined by the NHS (S1) (see Sup. Inf.) and Unhealthy otherwise. Self rep. healthy comprises individuals who self-reported health of 3 (“Good”) or less. HIV is an example of chronic infection while all the diseases can fall in the category of immune and blood-related conditions. While separation between overtly healthy and disease groups may be expected, the novelty of our proposed metric lies in its ability to provide a generalizable, quantitative, and interpretable single numerical score that scales across heterogeneous conditions and noisy datasets, offering a standardized baseline for predictive stratification. The Numerical Immune Score (NIS) shows a clear separation between healthy and disease groups on the same standardized scale, with healthy individuals clustering at higher scores and more confined distributions. Disease states exhibit broader shapes and hence, greater variability and declining scores reflecting increased immune dysregulation. Overlap across conditions highlights NIS’s sensitivity to graded immune perturbations rather than binary health–disease discrimination.
Table 5 shows examples of CCIS and its corresponding NCCIS values for a sample of real-life and artificial cases. The second column of the table shows the NIS, for purposes of comparison.
It is worth noting that these examples were generated using the NHS values for normal healthy subjects and based on the provisional assumption that the distance between the upper and normal ranges is twice the standard deviation.
The examples were taken from [29] as they were intended as very preliminary tests. We are aware that their source clearly states that they are meant only for teaching purposes. Here, they are used for purposes of illustration only, and the development of the NIS and NCCIS is not dependent on them.
In Table 4, we provide Cohen’s d values that suggest practical applicability of the index, making it clinically relevant for triaging purposes even at the level of individual classification and prediction.
@@
The method is shown to be valid for predicting immune profiles/trends (variation over time) observed in diverse patient groups when considering socio-clinical metadata parameters such as age and sex, thereby suggesting their clinical applicability for purposes such as triaging and fast screening and extensions to deviant clinical cases in prospective studies. In Fig. 1, it is shown that reported medical conditions produced a higher average NCCIS value, but also induced some anomalous clustering. Furthermore, the distribution of scores was even wider than that of the NIS. We therefore conclude that the NCCIS did enhance the insights already gleaned from the NIS and the CCIS.Separation of groups of self-reported data suggests that they could be separated from those that were not actually healthy at the time of testing.
The self-reported health information may not always correspond to objectively verified clinical status. Thus, the observation possibly reflects statistical discrepancies between self-reported and haematological patterns rather than confirmed diagnostic misclassification. Accordingly, while individuals who self-reported not being healthy generally exhibited blood profiles consistent with non-healthy patterns, this is suggested as a potential interpretation rather than clinical distinction. The number of individuals in each cohort is shown in Table 3. One possible explanation for the bimodality of some of these plots is the nature of the source data as self-reported, where respondents may or may not have been ill at the time of giving their answers, or from the application of the separating cutoff value from NIS.
@@
Statistical tests assessed the relationship among the three individualised scores, NIS, NCCIS, and CCIS, for various health conditions. Table 6 presents significant statistical associations between these scores.
The Kruskal-Wallis test, a non-parametric alternative to ANOVA was performed due to the ordinal nature of the CCIS index. The significant p-value (<<0.001) indicates that there are statistically significant differences in the distribution of the NIS, NCCIS, and CCIS scores in the conditions.
A Pearson correlation assessed the linear relationship between NIS and NCCIS. The strong positive correlation coefficient (0.89) with a highly significant p-value (<< 0.0001) determined from a two-tailed independent samples t-test suggests a robust linear association between these two scores.
Lastly, a Wilcoxon Signed-Rank test compared paired differences between NIS and NCCIS. The significant result (p = 0.018) indicates that there are significant differences between these paired scores and therefore each can capture different aspects of blood data.